# Effects of polar structure and moving ejecta on the dynamics of SGRB jets

### v1

@@ 3 Results
In this section, we describe the results of our simulations. We will focus on overall differences driven by the presence/absence of characteristic components (funnel and wind) in each simulation. The potential for these differences to lead to observational tests are discussed in Sect. 4.
@@ 3.1 Funnel and wind effects
A first general consideration is that we find that the presence of a funnel has the strongest impact on early jet evolution, while the wind mostly affects the large-scale behavior of the outflow. Consider, for example, the case in which the jet is in its early stage, at $\rm{t}=0.022~{}\mathrm{s}$, drilling through the different media (shown in the upper left panel of Figure 2). There is a marked difference between the left and right sides of the upper left panel (that is, the cases in which the medium does not have and has a funnel, respectively). The top and bottom parts, instead, are similar, showing that the presence or absence of a wind is barely noticeable at these times. Instead, as time progresses, the presence of a wind becomes more noticeable. For the case in which $\rm{t}=0.124~{}\mathrm{s}$, there is a clear difference between the top and bottom parts of the bottom right panel of Figure 2 (i.e., those with a static and expanding medium, respectively). The static case is characterized by extended angular wings that reach all the way to $90^{\circ}$. The presence of wind, instead (bottom part of the panel), produces hydrodynamical collimation at a large distance, resulting in a more focused outflow. We now examine each case in more detail and quantify the differences.
Jets evolving through a medium that has a funnel evolve faster $\approx 10\%$ faster) compared to the case in which there is no funnel. In addition, the funnel produces a larger jet opening angle. Meanwhile, jets that evolve through expanding media take more time to break out, and the jet has a narrower opening angle. In the breakout, the opening angle for the material with $\Gamma=2$ is $6^{\circ}$ is wider in the static medium with a funnel (F_nW) while $6^{\circ}$ narrower in the expanding wind without a funnel (F_nW), compared to the static medium without a funnel (nF_nW, for which $\theta\approx 29^{\circ}$). After the breakout, the combination of the funnel and the expanding wind produces a narrower opening angle of the jet head. The material with $\Gamma=2$, in the breakout (and final integration time), had $\theta\approx 29^{\circ}$ ($\theta\approx 75^{\circ}$) and $\theta\approx 33^{\circ}$ ($\theta\approx 90^{\circ}$) for models F_W and nF_nW, respectively.
@@
An analysis of the energy ratios across our models reveals that the turbulence within the cocoon consists of both kinetic and thermal components. As the jet evolves, substantial thermal energy conversion into kinetic energy occurs, particularly in the outer cocoon layers. Following the breakout, the cocoon transitions to being primarily kinetically dominated in both models, with regions of enhanced kinetic energy concentrated near the jet core. In addition, the fact that the cocoon is kinetically dominated agrees with the results from Hamidani & Ioka (2021).
@@ 3.2 Jet structure
Figure 3 presents the opening angles ($\theta$) of the fast-moving material in the jet head of our models as a function of time (normalized to their corresponding breakout time, $t-t_{\rm{bo}}$). Specifically, the lines in the figure show the largest angle from the jet axis at which the material moves with a certain speed ($\Gamma_{\infty}=2,5,10$). Different colors and line styles are used to identify the Lorentz factor considered and the possible presence of a funnel. The top panel shows the result for static media, while the bottom panel shows the results in a pre-accelerated medium. The vertical dashed gray line indicates the breakout time. The faster the jet material, the narrower the opening angles of the fast-moving material. For example, for the final integration time in our simulations, the opening angle for $\Gamma_{\infty}=10$ is approximately half that of $\Gamma_{\infty}=2$, independently of the characteristics of the medium. In addition, the opening angles of the jets are significantly narrower when propagating through an expanding medium. For example, the final opening angle in an expanding medium for $\Gamma_{\infty}=2$ is always below $\sim 75^{\circ}$ while it reaches $90^{\circ}$ in a static medium ($\Gamma_{\infty}=2,5$ are reduced by $\sim 10^{\circ}$ and $\sim 5^{\circ}$, respectively).
@@
To identify the angular energy distribution ($dE/d\Omega$) for the different models, we use the following equation:
@@ 4 Summary & discussion
In this section, we summarize and discuss our results in the context of other studies in the field. In addition, we discuss what aspects are amenable to be tested with current or future observations.
This study explores the influence of two properties of the ambient material surrounding a BNS merger on the propagation of a relativistic outflow. These two properties are polar stratification (namely, the presence of a low-density funnel along the angular momentum direction) and the role of preexisting outward velocity. Our simulations demonstrate that a static funnel geometry enhances lateral cocoon expansion, increases the forward shock distance, and broadens the cocoon structure, especially in the early phase of the jet evolving through the media. The effect of the outward motion of the medium, instead, manifests itself at later times when a more effective collimation of the relativistic material is evidenced. Since the two explored components have an impact in different phases, it is not surprising that they are additive. When both are present, the effect is that of a funnel-driven early evolution coupled with increased collimation at later stages.
@@
Our results for the expanding medium models show that the wind may affect the evolution of the jet and are consistent with the findings of Murguia-Berthier et al. (2017). However, we obtain angular energy values that are $\sim$ three orders of magnitude larger (we notice that their jets are less energetic and that their ambient medium is different from ours, making a direct comparison not straightforward). In agreement with Lazzati & Perna (2019), who investigated jets propagating through media with varying masses and velocities, we found that $\theta_{\rm{j}}$ and $\theta_{\rm{c}}$ are broader in lower-density environments. The expanding wind with a funnel model produces angular energy distributions similar to those reported by Geng et al. (2019), where the medium expands at $0.2\rm{c}$ and has a low-dense polar structure ($\rho\propto\sin^{3}\theta$). Our energy profile shows a well-defined jet followed by a plateau if the cocoon evolves through an expanding medium. In contrast, they obtain narrower jets (by $\sim 5^{\circ}$) and their absence of data beyond $\theta\geq 50^{\circ}$ limits a comparison of the shocked material at large viewing angles. Our findings align with those of Pais et al. (2024), which showed that a narrow funnel acts as a barrier leading to collimated outflows. In addition, we confirm that jets with high luminosities ($\sim 10^{51}$ erg s-1) lack recollimation shocks within the jet. To check that this is not a numerical effect or due to our numerical setup and geometry, we ran an additional simulation with lower jet luminosity ($5\times 10^{50}$ erg s-1). We confirmed that recollimation shocks appear in the lower luminosity jet case.
We note that the present study is constrained to two-dimensional axisymmetric simulations in a relatively small domain (at least one or two orders of magnitude short of the photospheric radius) and no radiation transport was included. The jet is taken to have a constant luminosity and its magnetization, as well as that of the medium, is not taken into account. In addition, the medium was modeled to expand with a constant velocity. Future investigations are required to explore the impact of these simplifications, including the extension to a larger three-dimensional computational domain, variable jet behavior, the role of magnetization, and the emission that would be produced.
@@ Acknowledgements
LGG and DLC acknowledge the support of the Miztli-UNAM supercomputer (project LANCAD-UNAM-DGTIC-321) and AWS (in the "Proyectos de investigación en la Nube UNAM-AWS") for the computational time assigned to perform the simulations and also acknowledge the support of the UNAM-PAPIIT grant IG100820. DL acknowledges support from the NSF grant AST-1907955.

### v2

@@ 3 Results
In this section, we describe the results of our simulations. We focus on overall differences driven by the presence or absence of characteristic components (funnel and wind) in each simulation. The implications of these differences for potential observational effects are discussed in Section 4.
@@ 3.1 Funnel and wind effects
A first general consideration that we find is that the presence of a funnel has the strongest impact on early jet evolution, while the wind mostly affects the large-scale behavior of the outflow. Consider, for example, the case in which the jet is in its early stage, at $\rm{t}=0.022~{}\mathrm{s}$, drilling through the different media (shown in the upper left panel of Figure 2). There is a marked difference between the left and right sides of the upper left panel (that is, the cases in which the medium does not have and has a funnel, respectively). The top and bottom parts, instead, are similar, showing that the presence or absence of a wind is barely noticeable at these times. Instead, as time progresses, the presence of a wind becomes more noticeable. For the case in which $\rm{t}=0.124~{}\mathrm{s}$, there is a clear difference between the top and bottom parts of the bottom right panel of Figure 2 (i.e., those with a static and expanding medium, respectively). The static case is characterized by extended angular wings that reach all the way to $90^{\circ}$. The presence of wind, instead (bottom part of the panel), produces hydrodynamical collimation at a large distance, resulting in a more focused outflow. We now examine each case in more detail and quantify the differences.
Jets evolving through a medium that has a funnel evolve faster $\approx 10\%$ faster) compared to the case in which there is no funnel. In addition, the funnel produces a larger jet opening angle. The latter two funnel effects are visible in Figure 7, where the evolution of the jet head over time for various models is shown. Meanwhile, jets that evolve through expanding media take more time to break out (see the dots and triangles in Figure 7).
In the breakout, the opening angle for the material with $\Gamma=2$ is $6^{\circ}$ is wider in the static medium with a funnel (F_nW) while $6^{\circ}$ narrower in the expanding wind without a funnel (F_nW), compared to the static medium without a funnel (nF_nW, for which $\theta\approx 29^{\circ}$). After the breakout, the combination of the funnel and the expanding wind produces a narrower opening angle of the jet head. The material with $\Gamma=2$, in the breakout (and final integration time), had $\theta\approx 29^{\circ}$ ($\theta\approx 75^{\circ}$) and $\theta\approx 33^{\circ}$ ($\theta\approx 90^{\circ}$) for models F_W and nF_nW, respectively.
@@
An analysis of the thermal and kinetic energy reveals that the material within the cocoon consists of both components (see for example Figure 8 where the spatial and temporal evolution of the energy ratio, $e_{k}/e_{th}$, for various models is shown). As the jet evolves, substantial thermal energy conversion into kinetic energy occurs, particularly in the outer cocoon layers. Following the breakout, the cocoon transitions to being primarily kinetically dominated in both models, with regions of enhanced kinetic energy concentrated near the jet core. In addition, the fact that the cocoon is kinetically dominated agrees with the results from Hamidani & Ioka (2021).
@@ 3.2 Jet structure
Figure 3 presents the opening angles of the fast-moving material in the jet-cocoon of our models as a function of time (normalized to their corresponding breakout time, $t-t_{\rm{bo}}$). The opening angle of material moving at a certain Lorentz factor $\Gamma$ is defined as follows. We first select all the parcels of material that move faster than the selected value of $\Gamma$. For each of these parcels we know their off-axis angle as their polar coordinate. We select the one with the largest value of polar angle and define that angle as the angle $\theta$ that is plotted in Figure 3. Different colors and line styles are used to identify the Lorentz factor considered and the possible presence of a funnel. The top panel shows the result for static media, while the bottom panel shows the results in a pre-accelerated medium. The vertical dashed gray line indicates the breakout time. The faster the jet material, the narrower the opening angles of the fast-moving material. For example, for the final integration time in our simulations, the opening angle for $\Gamma_{\infty}=10$ is approximately half that of $\Gamma_{\infty}=2$, independently of the characteristics of the medium. In addition, the opening angles of the jets are significantly narrower when propagating through an expanding medium. For example, the final opening angle in an expanding medium for $\Gamma_{\infty}=2$ is always below $\sim 75^{\circ}$ while it reaches $90^{\circ}$ in a static medium ($\Gamma_{\infty}=2,5$ are reduced by $\sim 10^{\circ}$ and $\sim 5^{\circ}$, respectively).
@@
[…]
@@ 4 Summary & discussion
In this section, we summarize and discuss our results in the context of other studies in the field. In addition, we discuss what aspects are amenable to being tested with current or future observations.
This study explores the influence of two properties of the ambient material surrounding a BNS merger on the propagation of a relativistic outflow. These two properties are polar stratification (namely, the presence of a low-density funnel along the angular momentum direction) and the role of preexisting outward velocity. Our simulations demonstrate that a static funnel geometry enhances lateral cocoon expansion, increases the forward shock distance, and broadens the cocoon structure, especially in the early phase of the jet evolving through the media. The effect of the outward motion of the medium, instead, manifests itself at later times when a more effective collimation of the relativistic material is evidenced. Since the two explored components have an impact in different phases, it is not surprising that they are additive. When both are present, the effect is that of a funnel-driven early evolution coupled with increased collimation at later stages. Additionally, the jets are collimated within the cocoon as the collimation condition, $\tilde{L}<\theta_{0}^{-4/3}$ (where $\tilde{L}=\rho_{j}h_{j}\Gamma_{j}/\rho_{f}$ and $\theta_{j}^{-4/3}$) of Bromberg et al. (2011), is satisfied. Specifically, for the values in our models, we have $\tilde{L}\sim 2$ and $\theta_{j}^{-4/3}\sim 25$).
@@
Our results for the expanding medium models show that the wind may affect the evolution of the jet and are consistent with the findings of Aloy et al. (2005), who demonstrated that low-density polar funnels enhance jet acceleration, while Murguia-Berthier et al. (2017) showed that winds can restrict lateral jet expansion. However, we obtain angular energy values that are $\sim$ three orders of magnitude larger (we notice that their jets are less energetic and that their ambient medium is different from ours, making a direct comparison not straightforward). Additionally, Nativi et al. (2021) further showed that asymmetric winds can effectively collimate the jet. In agreement with Lazzati & Perna (2019), who investigated jets propagating through media with varying masses and velocities, we found that $\theta_{\rm{j}}$ and $\theta_{\rm{c}}$ are broader in lower-density environments. The expanding wind with a funnel model produces angular energy distributions similar to those reported by Geng et al. (2019), where the medium expands at $0.2\rm{c}$ and has a low-density polar structure ($\rho\propto\sin^{3}\theta$). Our energy profile shows a well-defined jet followed by a plateau if the cocoon evolves through an expanding medium. In contrast, they obtain narrower jets (by $\sim 5^{\circ}$) and their absence of data beyond $\theta\geq 50^{\circ}$ limits a comparison of the shocked material at large viewing angles.
Our results are consistent with those of Pavan et al. (2021), whose analysis demonstrates that a low-density funnel enhances jet breakout efficiency, enabling the jet to retain higher energy by reducing baryon loading and interaction losses during its propagation through the ejecta. This is consistent with our results, where we also observe that the presence of a low-density funnel prevents jet lateral expansion and boosts jet velocity. Our findings align with those of Pais et al. (2024), which showed that a narrow funnel acts as a barrier leading to collimated outflows. In addition, we confirm that jets with high luminosities ($\sim 10^{51}$ erg s-1) lack recollimation shocks within the jet. To check that this is not a numerical effect or due to our numerical setup and geometry, we ran an additional simulation with lower jet luminosity ($5\times 10^{50}$ erg s-1). We confirmed that recollimation shocks appear in the lower luminosity jet case.
We note that the present study is constrained to two-dimensional axisymmetric RHD simulations in a relatively small domain (at least one or two orders of magnitude short of the photospheric radius) and no radiation transport was included. The jet is taken to have a constant luminosity and its magnetization, as well as that of the medium, is not taken into account. In addition, the medium was modeled to expand with a constant velocity. Future investigations are required to explore the impact of these simplifications, including the extension to a larger three-dimensional computational domain, variable jet behavior, the role of magnetization, and the emission that would be produced.
@@ Acknowledgements
We thank Fabio De Colle and Enrique Moreno Méndez for their valuable comments and suggestions which helped improve the quality of this work. LGG and DLC acknowledge the support of the Miztli-UNAM supercomputer (project LANCAD-UNAM-DGTIC-321) and AWS (in the "Proyectos de investigación en la Nube UNAM-AWS") for the computational time assigned to perform the simulations and also acknowledge the support of the UNAM-PAPIIT grant IG100820. DL acknowledges support from the NSF grant AST-1907955.
@@
Appendix A Initial Density Distribution
The initial density profile for the model F_nW. The density is modeled using Equation 1 with $\eta=6$ and $\theta_{f}=30^{\circ}$. The color map represents the logarithmic scale of the density $\log_{10}\rho$ in units of $\mathrm{g,cm^{-3}}$. Iso-density contours are overlaid on the plot, shown in black, to highlight the levels of constant density.
Figure 6: Initial density profile for model F_nW (model with $\eta=6$ and $\theta_{f}=30^{\circ}$). The 2D color map shows $\log_{10}\rho$ in units of $\mathrm{g\,cm^{-3}}$. The black contours indicate iso-density levels ($\rho=(1,10^{2},10^{4},10^{6})\mathrm{g\,cm^{-3}}$).
Appendix B Jet-Head
Figure 7 presents the evolution of the jet head over time for various models (F_nW: solid red line, F_W: solid purple line, nF_W: dashed purple line, and nF_nW: dashed red line). Dots and triangles mark the breakout time for each model: 0.045 s for F_nW, 0.05625 s for F_W, 0.0525 s for nF_W, and 0.0525 s for nF_nW. The curves are truncated shortly after breakout, as all models continue to follow the same overall trend beyond this point.
Figure 7: Evolution of the jet head over time for various models (F_nW: solid red line, F_W: solid purple line, nF_W: dashed purple line, and nF_nW: dashed red line). Dots and triangles represent the breakout time for each model.
Appendix C Thermal to kinetic energy
Figure 8 presents kinetic-to-thermal energy ratio ($e_{k}/e_{th}$) for a relativistic jet moving through a static medium without funnel (nF_nW, upper left), the static medium with funnel (F_nW, upper right), the expanding medium without funnel (nF_W, lower left). The evolution of the jet is presented at four different times $\rm{t}=0.022~{}\mathrm{s}$, $0.045~{}\mathrm{s}$, $0.086~{}\mathrm{s}$, and $0.124~{}\mathrm{s}$.
Figure 8: Kinetic-to-thermal energy ratio ($e_{k}/e_{th}$) for a relativistic jet moving through a static medium without funnel (nF_nW, upper left), the static medium with funnel (F_nW, upper right), the expanding medium without funnel (nF_W, lower left), and the expanding medium with funnel (F_W, lower right). Four different times are shown ($\rm{t}=0.022~{}\mathrm{s}$, $0.045~{}\mathrm{s}$, $0.086~{}\mathrm{s}$, and $0.124~{}\mathrm{s}$). Note that the spatial domain sizes differ between the upper and lower panels.